OpenStudy (anonymous):

.Fifty students in the fourth grade class listed their hair and eye colors in the table below: Brown hair Blonde hair Total Green eyes 16 12 28 Brown eyes 14 8 22 30 20 50 Are the events "green eyes" and "brown hair" independent? No, P(brown hair) • P(green eyes) ≠ P(brown hair ∩ green eyes Yes, P(brown hair) • P(green eyes) ≠ P(brown hair ∩ green eyes) No, P(brown hair) • P(green eyes) = P(brown hair ∩ green eyes) Yes, P(brown hair) • P(green eyes) = P(brown hair ∩ green eyes

OpenStudy (anonymous):

@1lorax2

OpenStudy (amistre64):

how do we define independence?

OpenStudy (anonymous):

i dont have a clue ;-;

OpenStudy (amistre64):

independence is yes, if they are equal; meaning that one has no effect on the other it is no if they are not equal.

OpenStudy (anonymous):

So, im going with a Yes answer?

OpenStudy (amistre64):

i dont know if they are equal or not :/ what do you get for the probs?

OpenStudy (anonymous):



OpenStudy (anonymous):

I Know the chart looked a little hard to read xD

OpenStudy (amistre64):

P(BrH) = 30/50 P(GrE) = 28/50 P(BrnGr) = 16/50 do these probs look right?

OpenStudy (anonymous):

Yess

OpenStudy (amistre64):

then, do they equal, or not equal when used?

OpenStudy (anonymous):

Equal?

OpenStudy (amistre64):

guessing doesnt count .. another way is to compare rows/cols for equal probability: does 16/28 = 30/50? \[\frac{30(28)}{50(50)}=\frac{16}{50}\] \[\frac{30}{50}=\frac{16}{28}\]

OpenStudy (anonymous):

Nope, it doesnt.

OpenStudy (amistre64):

then id go with no, since they are not equal

OpenStudy (anonymous):

Soo, im going with A?

OpenStudy (amistre64):

there is no A in the options, but im assuming you mean the first one

OpenStudy (anonymous):

Yeah, the first one.

OpenStudy (amistre64):

then yeah :)

OpenStudy (anonymous):

Thaanks ;o